OpenStudy (osanseviero):

g(4)=-14 g'(4)=-5 g(3)=-18 g'(3)=-3 The equation is probably a quadratic function, but is there a way to get it?

OpenStudy (loser66):

to me, I will do something like this the quadratic function has the form g(x) = ax^2 +bx +c g(2) = 4a + 2b +c = -14 g(3) = 9a +3b +c = -18

OpenStudy (loser66):

now, take derivative: g'(x) = 2ax +b g'(2) = 4a + b = -5 g'(3) = 6a +b = -3 so that 2a = 2 therefore a =1 and b = -9

OpenStudy (osanseviero):

oh...thanks!

OpenStudy (loser66):

you got it? still have to find c

OpenStudy (osanseviero):

Yep, but it is easy to find now :) thanks :) got it perfectly

OpenStudy (osanseviero):

c=0

OpenStudy (osanseviero):

Perfect :) thanks a lot

OpenStudy (loser66):

:)

OpenStudy (osanseviero):

I got a problem related to this...do you think you could help me?

OpenStudy (loser66):

not sure, if it is not over my head, I will